OpenStudy (anonymous):

t^2+6=8t solve the equation by completing the square and leave the irrational roots in simplest radical form. please help solve.

OpenStudy (anonymous):

just like the last one but this time start with \[t^2-8t=-6\]

OpenStudy (anonymous):

satellite please help meee

OpenStudy (anonymous):

you get \[(t-4)^2=-6+4^2=-6+16=10\] \[t-4=\pm\sqrt{10}\] \[t=4\pm\sqrt{10}\]

OpenStudy (anonymous):

thank you satellite. you are really good in math!

OpenStudy (anonymous):

yw